a person standing on the top of a hill 30m high has to throw a ball to his frind standing on the ground 40m horizontally away.if he throws the all directly aminig at his friend with  speed of 125/3 packet fallls at a distance of 20/am from hids friend .a is an integer then find value of a 

Grade:11

1 Answers

Piyush
8 years ago
Apply
x= uxt +(1/2)axt2 where ax=0
Now to calculate t Apply
y=(1/2)gt2 where uy=0
substitute the value of x and y you will get ‘a’.